# Quantifying Fire Risk Index in Chemical Industry Using Statistical Modeling Procedure

### v1

@@ 3.3 Cluster Interpretation
Cluster 10 Words 1 distillation column, exposure, solvent, photoinitiator, metal, cleaning solvent, heptane, nozzle, flexible, silanes manufacturing 2 ignition source, polishing, painting, floor, laboratory, electricity, plenty of, scrap paper, waste, recycling 3 incinerate, waste wood, interior material, heat, urethane, smoldering ignition, cutting machine, dust collection equipment, rubber, manufacturing machine 4 heat of reaction, expired reagent, corn, cooking oil, storage tank, reagent, oxygen, eruption, heat wave, rainwater 5 abnormal reaction, pharmaceutical, film, sheath heater, upper, suction, oil, wire mesh, epoxy, dust explosion 6 storage, cable, malfunction, cool down, oil tank, control box, coil, electrical circuit board, refrigerator, fertilizer 7 coating, injection, corrosion, cutting oil, grinder, press, repair, acetylene, scrubber, steam equipment 8 vapor, nucleic acid, grease, leak, hazardous material, mix, paint, high temperature, large-scale, thermal cutting 9 drying room, cosmetic, pallet, ventilation fan, pressure, ceiling, rotation, small amount, decompose, lid 10 preheat, ignition, stacked, accumulation, tire, moisture, activated carbon, semi-finished product, waste storage, welding 11 cooling fan, service line, shut out, lower terminal, insulating oil, underground, fan belt, switch, electric current, analysis lab 12 closing, charging equipment, flammability, maintenance, automation, repair, cooling machine, machine room, gas torch, arc 13 dilution, power outage, cover, immediate upper, tracking, prefabricated, vacuum, muller, vulcanizer, melting fusion 14 micro, seat pad, deodorizing tower, absorbent pad, gunnysack, ton bag, compost, boiler room, sesame dregs, oxidation heat 15 belt, particle, capture, aluminum, inlet, rubber department, wood chip, blowing, metal powder, abradant
@@
Taken together, the interpretation of clusters provides meaningful categorical insights into the semantic structure of fire investigation records. However, semantic similarity alone does not capture the extent of economic severity associated with each term. To address this, we incorporated fire-damage estimates into the analysis by estimating word-level coefficients via LASSO regression. This approach allows us to link clusters with the magnitude of potential financial losses. The coefficients estimated for each word quantify its relative contribution to explaining variations in property damage estimates. Figure 3 presents a three-dimensional visualization, where the horizontal axes represent the latent positions of words and the vertical axis corresponds to their regression coefficients. The clusters are indicated by color, consistent with the grouping shown in Figure 2.
@@
Figure 3: Three-dimensional visualization of words, where the $x$-and $y$-axes represent latent positions and the $z$-axis corresponds to coefficients estmated from a LASSO regression on fire-damage insurance claims. Colors indicate clusters identified through $k$-means grouping.
@@ 3.4 Risk Index Estimation
The index $\gamma_{i,c}$ measures the relative contribution of word $i$ within cluster $c$ to the estimation of property damage amounts. A higher value indicates that the word is more strongly associated with larger expected losses compared to other words in the same cluster. Table 4 presents the top ten words with the highest $\gamma_{i,c}$ values in each cluster, along with their Risk Index. For example, in Cluster 2 (Ignition from heat accumulation near combustible interior materials) and Cluster 7 (Ignition of flammable substances by friction heat or spark from machinery), words such as flame, heat, and grinder show high $\gamma_{i,c}$ values, reflecting their strong linkage with higher levels of estimated property damage. Thus, $\gamma_{i,c}$ highlights words whose relative importance provides insight into the financial risk implications captured within each cluster.
@@ 4 Discussion
We instantiate this framework via the Overall Word Risk Index, $\rho_{i,c}$, which quantifies the association between terms and loss outcomes rather than raw frequency. Using $\rho_{i,c}$ to rank and organize the vocabulary in the embedding (Fig. 4), we identify four cross-cutting facets of fire risk: (i) chemical leakage/vapors (red), (ii) storage/warehouses (blue), (iii) equipment/electrical (green), and (iv) self-ignition (pink). In what follows, we analyze each facet and its representative cases.
@@
Second, structural and operational factors such as warehouses, storage facilities, and pallets are also prominent, underscoring the critical role of storage conditions and facility maintenance. Prior research on cold-chain logistics demonstrates that pallet stacking and package design strongly influence airflow and heat transfer within storage rooms, where insufficient ventilation can generate localized hotspots and elevate fire hazards [1]. In the context of energy storage, investigations highlight that inadequate cooling and airtight designs in portable energy storage systems promote overheating and exacerbate ignition risk [6]. Similarly, studies on lithium-ion batteries reveal that poor thermal management and accumulation of flammable gases in confined storage spaces can accelerate thermal runaway and combustion [7]. Together, these findings indicate that storage practices—ranging from palletized materials to battery warehouses—interact with flammable substances in ways that compound ignition potential, while recurring references to ventilation-related equipment emphasize how failures in airflow management can amplify both the likelihood and severity of fire incidents.
@@
Although cluster membership (shapes in Fig. 4) reflects lexical neighborhoods, nearby terms can encode different operational contexts that co-occur in narratives. For example, storage terms (blue) can sit beside equipment/electrical terms (green) because pallet stacking and enclosure geometry restrict ventilation and heat rejection [1], thereby increasing the efficacy of routine electrical faults as ignition sources; likewise, self-ignition terms (pink) may appear near leakage/vapor terms (red) when oxygen ingress and heat accumulation in stacked goods elevate vapor formation and ignition potential [23]. Conversely, process-centric terms (e.g., maintenance/automation) may be spatially offset when hot-work contexts (sparks, localized heating) are discussed apart from storage constraints, yet they remain semantically bridgeable whenever operations occur proximate to flammable inventories or confined airflow paths [5, 7]. Interpreting the embedding jointly with the loss-aware index $\rho_{i,c}$ (Fig. 4) discriminates semantically adjacent yet economically distinct patterns: terms that lie close in the map but carry higher $\rho_{i,c}$ mark contexts historically associated with larger property losses, whereas nearby low-$\rho_{i,c}$ terms indicate operational exposure without the same tail severity. This joint reading—proximity for mechanism, $\rho_{i,c}$ for consequence—yields a ranked set of actionable priorities, directing inspection toward leakage/vapor configurations while maintaining vigilance for storage and self-ignition scenarios that can escalate under adverse conditions.
Methodologically, our analysis proceeds in four stages. First, we transform unstructured investigation narratives into topical structure via topic modeling and latent embeddings, which provide low-dimensional coordinates for terms and documents. Second, we quantify word–word interactions by combining co-occurrence statistics with embedding-based proximity to form a network. Third, we partition this network into semantically coherent clusters and assign interpretable labels using representative cases to recover operational and environmental contexts. Fourth, we couple these text-derived signals with structured loss data variables to construct a loss-aware score—the Overall Word Risk Index $\rho_{i,c}$—that prioritizes terms by their association with financial outcomes. Interpreting the embedding jointly with $\rho_{i,c}$ thus yields a map that is both mechanistic (pathways reflected in neighborhoods) and consequential (loss-linked salience), enabling prevention, inspection, and maintenance efforts to focus on the most actionable risks.

### v2

@@ 3.2 Thematic Aggregation of Topics
Cluster 10 Words 1 distillation column, exposure, solvent, photoinitiator, metal, cleaning solvent, heptane, nozzle, flexible, silanes manufacturing 2 ignition source, polishing, painting, floor, laboratory, electricity, plenty of, scrap paper, waste, recycling 3 incinerate, waste wood, interior material, heat, urethane, smoldering ignition, cutting machine, dust collection equipment, rubber, manufacturing machine 4 heat of reaction, expired reagent, corn, cooking oil, storage tank, reagent, oxygen, eruption, heat wave, rainwater 5 abnormal reaction, pharmaceutical, film, sheath heater, upper, suction, oil, wire mesh, epoxy, dust explosion 6 storage, cable, malfunction, cool down, oil tank, control box, coil, electrical circuit board, refrigerator, fertilizer 7 coating, injection, corrosion, cutting oil, grinder, press, repair, acetylene, scrubber, steam equipment 8 vapor, nucleic acid, grease, leak, hazardous material, mix, paint, high temperature, large scale, thermal cutting 9 drying room, cosmetic, pallet, ventilation fan, pressure, ceiling, rotation, small amount, decompose, lid 10 preheat, ignition, stacked, accumulation, tire, moisture, activated carbon, semi-finished product, waste storage, welding 11 cooling fan, service line, shut out, lower terminal, insulating oil, underground, fan belt, switch, electric current, analysis lab 12 closing, charging equipment, flammability, maintenance, automation, repair, cooling machine, machine room, gas torch, arc 13 dilution, power outage, cover, immediate upper, tracking, prefabricated, vacuum, muller, vulcanizer, melting fusion 14 micro, seat pad, deodorizing tower, absorbent pad, gunnysack, ton bag, compost, boiler room, sesame dregs, oxidation heat 15 belt, particle, capture, aluminum, inlet, rubber department, wood chip, blowing, metal powder, abradant
@@
Taken together, the interpretation of clusters provides meaningful categorical insights into the semantic structure of fire investigation records. However, semantic similarity alone does not capture the extent of economic severity associated with each term. To address this, we incorporated fire-damage estimates into the analysis by estimating word-level coefficients via LASSO regression. This approach allows us to link clusters with the magnitude of potential financial losses. The coefficients estimated for each word quantify its relative contribution to explaining variations in property damage estimates. Figure 3 presents a three-dimensional visualization, where the horizontal axes represent the latent positions of words and the vertical axis corresponds to their regression coefficients. Each colored dot denotes a single word, and the color indicates its cluster membership consistent with the grouping shown in Figure 2. This visualization allows direct comparison between semantic proximity (latent positions) and the estimated contribution of each word to financial damage. Words located at higher positions along the coefficient axis correspond to terms associated with higher risk of financial loss, whereas clusters with generally lower coefficient values indicate groups of words linked to less severe incidents. By examining the clusters from multiple viewing angles, one can observe that several clusters exhibit a wide range of coefficient magnitudes, suggesting heterogeneous risk levels within the same semantic domain, while others maintain consistently low coefficients, reflecting more homogeneous, low-risk contexts.
@@
Figure 3: Three-dimensional visualization of latent word embeddings from multiple perspectives. The $x$- and $y$-axes represent latent positions, while the $z$-axis corresponds to coefficients estimated from the LASSO regression on fire-damage insurance claims. Panels (a), (b), and (c) show the front, side, and top views, respectively. Colors indicate clusters identified through k-means grouping.
@@ 3.3 Risk Index Estimation
As described in Step 4 of Figure 1, the Lasso-based risk modeling was applied to estimate the $\gamma_{i,c}$ that measures the relative contribution of word $i$ within cluster $c$ to the estimation of property damage amounts. A higher value indicates that the word is more strongly associated with larger expected losses compared to other words in the same cluster. Table 5 presents the top ten words with the highest $\gamma_{i,c}$ values in each cluster, along with their Risk Index. For example, in Cluster 2 (Ignition from heat accumulation near combustible interior materials) and Cluster 7 (Ignition of flammable substances by friction heat or spark from machinery), words such as flame, heat, and grinder show high $\gamma_{i,c}$ values, reflecting their strong linkage with higher levels of estimated property damage. Thus, $\gamma_{i,c}$ highlights words whose relative importance provides insight into the financial risk implications captured within each cluster.
@@ 4 Discussion
We instantiate this framework via the Overall Word Risk Index, $\rho_{i,c}$, which quantifies the association between terms and loss outcomes rather than raw frequency. Using $\rho_{i,c}$ to rank and organize the vocabulary in the embedding (Figure 4), we identify four cross-cutting facets of fire risk: (i) chemical leakage/vapors (red), (ii) storage/warehouses (blue), (iii) equipment/electrical (green), and (iv) self-ignition (pink). The latent coordinates in Figures 2 and 4 are derived from the same embedding space, while Figure 4 highlights representative high-risk keywords and their corresponding topics to provide a more detailed interpretation across the four facets. In what follows, we analyze each facet and its representative cases.
@@
Second, structural and operational factors such as warehouses, storage facilities, and pallets are also prominent, underscoring the critical role of storage conditions and facility maintenance. Prior research on cold-chain logistics demonstrates that pallet stacking and package design strongly influence airflow and heat transfer within storage rooms, where insufficient ventilation can generate localized hotspots and elevate fire hazards [1]. In the context of energy storage, investigations highlight that inadequate cooling and airtight designs in portable energy storage systems promote overheating and exacerbate ignition risk [10]. Similarly, studies on lithium-ion batteries reveal that poor thermal management and accumulation of flammable gases in confined storage spaces can accelerate thermal runaway and combustion [11]. Together, these findings indicate that storage practices, ranging from palletized materials to battery warehouses, interact with flammable substances in ways that compound ignition potential, while recurring references to ventilation-related equipment emphasize how failures in airflow management can amplify both the likelihood and severity of fire incidents.
@@
Although cluster membership (shapes in Fig. 4) reflects lexical neighborhoods, nearby terms can encode different operational contexts that co-occur in narratives. For example, storage terms (blue) can sit beside equipment/electrical terms (green) because pallet stacking and enclosure geometry restrict ventilation and heat rejection [1], thereby increasing the efficacy of routine electrical faults as ignition sources; likewise, self-ignition terms (pink) may appear near leakage/vapor terms (red) when oxygen ingress and heat accumulation in stacked goods elevate vapor formation and ignition potential [39]. Conversely, process-centric terms (e.g., maintenance/automation) may be spatially offset when hot-work contexts (sparks, localized heating) are discussed apart from storage constraints, yet they remain semantically bridgeable whenever operations occur proximate to flammable inventories or confined airflow paths [9, 11]. Interpreting the embedding jointly with the loss-aware index $\rho_{i,c}$ (Fig. 4) discriminates semantically adjacent yet economically distinct patterns: terms that lie close in the map but carry higher $\rho_{i,c}$ mark contexts historically associated with larger property losses, whereas nearby low-$\rho_{i,c}$ terms indicate operational exposure without the same tail severity. This joint reading proximity for mechanism, $\rho_{i,c}$ for consequence which yields a ranked set of actionable priorities, directing inspection toward leakage/vapor configurations while maintaining vigilance for storage and self-ignition scenarios that can escalate under adverse conditions.
The proposed framework not only enables post-incident classification of fire causes but also provides a practical foundation for preventive fire management. By translating text-derived risk indicators into measurable patterns, the derived Risk Index helps identify high-risk operational contexts and materials before incidents occur. Table 8 illustrates how the ranked Risk Index can guide on-site inspections and safety checklists in industrial facilities. This approach supports decision-makers in prioritizing safety measures under limited manpower and budget, and establishes a systematic linkage between historical fire evidence and actionable preventive guidelines. Moreover, the framework can be extended by incorporating real-time sensor-based parameters such as temperature, dust concentration, or gas-leakage levels to enable proactive monitoring analogous to diagnostic systems in medicine. In practice, these parameters can be included as additional covariates in the Risk Index estimation stage, allowing the model to quantify how environmental conditions influence fire risk levels. By jointly modeling textual indicators and measurable physical factors, the extended framework can provide more timely and context-aware assessments, supporting preventive decision-making in high-risk industrial settings.
Table 8: Example of Fire Prevention Guidelines for Industrial Facilities by Risk Index Ranking
Risk Index Cluster No. Topic Main Keywords Related Words Fire and Explosion Prevention Checklist 1.000 8 Ignition of Flammable Vapor during Organic Solvent Use chemical material, hazardous material leak, paint, sludge, grease, vapor, mixing equipment Are measures in place to prevent leakage of flammable chemical materials? Is there any accumulation of sludge such as grease? Is the ventilation system operating to prevent retention of flammable vapors? Are vapors accumulating where sparks from mixing equipment could ignite? 0.977 6 Electrical Fires in Areas Handling Flammable Materials shipping area, distribution board Are wiring systems, including distribution boards in the shipping area, properly maintained? 0.866 9 Ignition from Temperature Rise in Oil Vapor Areas warehouse, breakdown, discharge, storage, secondary battery, synthetic resins, hot work equipment secondary battery, discharge, hot work equipment, synthetic resins Are discharged secondary batteries stored safely? Is hot work inside the warehouse performed under safe conditions? Are combustible materials stored at a safe distance from hot work areas?
5 Conclusion
This study presented a data-driven framework for quantifying and interpreting fire-related risk factors embedded in unstructured investigation narratives. By transforming textual records into latent topical structures and constructing an embedding-based network of word interactions, we identified semantically coherent clusters that capture operational and environmental contexts of fire incidents. The key contribution of this work lies in linking these text-derived semantic signals with structured financial loss data to develop a composite metric, the Overall Word Risk Index ($\rho_{i,c}$), which highlights words and topics most closely associated with real-world damage outcomes. The resulting framework provides an interpretable and loss-aware map of industrial fire risk, offering a foundation for evidence-based prevention and safety management.
In this study, the derived Risk Index can be considered useful as reference data for safety management and insurance underwriting of industrial sites. Within a workplace, the Risk Index can help assign priorities when establishing preventive measures against potential accidents inferred from topics and words. Given the limited manpower and budget typically available at industrial sites, allocating safety management resources in proportion to the degree of risk is of critical importance. In the context of insurance subscription, understanding the level of risk at a workplace is also of utmost priority. Although insurance premiums are generally determined based on statistical loss ratios by industrial sector, it remains difficult to assess the unique risks arising from a site’s specific operating conditions. In this regard, the Risk Index serves as a quantitative measure of a facility’s risk level and may provide valuable guidance for insurance underwriters when determining whether to underwrite facilities that involve substances or processes associated with higher risk scores.
Although the results are promising, potential threats to validity should also be acknowledged. The narratives used in this study were prepared by professional fire officers based on verified evidence and standardized terminology, ensuring a high degree of internal validity with minimal subjective bias. While supervised validation would ideally complement our approach, the current unsupervised framework incorporates expert interpretation to maintain analytical reliability. Regarding external validity, the proposed methodology could be extended to other domains such as manufacturing, logistics, construction, or residential and public facilities to assess its generalizability and adaptability across diverse fire contexts.